# Feature Denoising for Low-Light Instance Segmentation using Weighted Non-Local Blocks

### v1

@@ Abstract
Instance segmentation for low-light imagery remains largely unexplored due to the challenges imposed by such conditions, for example shot noise due to low photon count, color distortions and reduced contrast. In this paper, we propose an end-to-end solution to address this challenging task. Based on Mask R-CNN, our proposed method implements weighted non-local (NL) blocks in the feature extractor. This integration enables an inherent denoising process at the feature level. As a result, our method eliminates the need for aligned ground truth images during training, thus supporting training on real-world low-light datasets. We introduce additional learnable weights at each layer in order to enhance the network’s adaptability to real-world noise characteristics, which affect different feature scales in different ways. Experimental results show that the proposed method outperforms the pretrained Mask R-CNN with an Average Precision (AP) improvement of +10.0, with the introduction of weighted NL Blocks further enhancing AP by +1.0.
@@
Index Terms— Instance segmentation, low-light images, feature denoising, non-local means, end-to-end
1 Introduction
@@
Instance segmentation is a crucial task in computer vision because it allows systems to interpret scenes with higher levels of detail, by understanding the position and shape of the objects within them. This can be highly beneficial in scenarios such as autonomous driving, video surveillance, augmented reality, multimedia editing, medical imaging and agricultural monitoring. Its importance is underscored by numerous recent research contributions in this field and various benchmarks and datasets, including Microsoft COCO [1], ADE20K [2], Pascal VOC [3] and Cityscapes [4].
@@
While its applications span many sectors, there has been limited research on instance segmentation in low-light conditions. The is due to the fact that low-light images are heavily affected by noise from a low photon count in darker conditions, making the task much more challenging. A common approach is to first apply a low-light image enhancement (LLIE) method to these images. This enables existing instance segmentation methods, typically designed for normal-light conditions, to run inference upon the enhanced images. LLIE methods include EnlightenGAN [7], Zero-DCE [8], and more recently RetinexFormer [6]. Specifically, Zero-DCE [8], and RetinexFormer [6] also use low-light object detection benchmarks (ExDark [9], DARK FACE [10]) to show their low-light enhancing performances.
To the best of our knowledge to date, the only work published on low light instance segmentation is by by Chen et al. [11]; however their method was only applicable for RAW images to take advantage of the higher bit-depth. In this paper, we propose an end-to-end approach that takes low-light sRGB images as inputs; these are much more widely used than RAW images and can be applied to video sequences.
@@
Our goal is to develop a robust and effective end-to-end method for low-light instance segmentation. This implies eliminating any explicit pre-processing steps in order to simplify the mask generating procedure. However most models are not designed to handle images with noise and low-contrast. In our experiments, we have found that both low-contrast and noise impact the performance of these models, as also confirmed by the study in [12].
Our focus is to build upon existing feature extractors to learn to ignore the presence of noise in the feature maps. This is achieved by integrating denoising blocks with non-local (NL) means as its denoising operation into the backbone. Specifically, our framework is based on Mask R-CNN [5] and leverages an NL block [13] for denoising in the feature extractor. We however extend this with a weighted NL block, where the additional learnable weights ensure that the network can learn the noise characteristics that affect each scale of feature space differently. We trained our framework with a synthetic dataset; this is due to the infancy of performing high-level computer vision tasks under low-light conditions. We used the popular Microsoft COCO [1] dataset and synthesised low-light conditions following the pipeline proposed by [14].
2 Related Work
2.1 Instance Segmentation
@@
Instance segmentation aims to produce a segmentation mask, bounding box and class label for every object detected in an image. One of the most popular techniques for achieving this is Mask R-CNN [5], introduced by He et al. as an extension to the popular object detection model, Faster R-CNN [15]. The architecture is the same except Mask R-CNN [5] introduced a new mask head to also predict the instance masks from the Regions of Interest (RoIs) proposed by the Region Proposal Network (RPN). With the notable work from [16], there has been a surge in transformer-based methods in recent years. For example, DETR-based [17] instance segmentation methods use CNN-based feature extraction backbones and concatenate positional embeddings before passing into a transformer.
2.2 Low-light image enhancement
@@
Traditional LLIE methods treat the problem as two separate tasks: contrast enhancement and denoising. Contrast enhancement methods include gamma correction and adaptive histogram equalizatiom (AHE) [18]; however these lack consideration of illumination factors, leading to cognition approaches inspired by Retinex Theory [19]. Contrast-enhancement of low-light images however, amplifies noise which necessitates the use of denoising algorithms for further refinement.
@@
Recent advancements in LLIE have explored the use of deep learning. Lore et al. first introduced LLNet [20], an autoencoder trained with synthetic low-light images. Supervised methods, including RetinexNet [21], KiND [22], and AGLLNet [14], require large sets of paired normal/low-light images for training, often generated synthetically. Unsupervised methods gained popularity due to the scarcity of paired normal-light images in low-light datasets. Examples include EnlightenGAN [7] and Zero-DCE (and its extended version Zero-DCE++) [8]. RetinexFormer [6], a transformer-based method, is currently the state-of-the-art approach, leveraging the Retinex theory [19] to separate color vision into reflectance and luminance components.
2.3 Low-light scene understanding
Several datasets have been released for higher computer vision tasks in low-light conditions. Semantic segmentation is a more common task, with datasets like NightCity [23] and BDD100K-night [24] focusing on analyzing road scenes to enhance autonomous driving at night. Additionally datasets such as ExDark [9] and DARK FACE [10] serve as benchmarks for low-light object detection.
Chen et al. [11] investigated instance segmentation in extreme low-light conditions, leveraging the higher bit-depth of RAW images to enhance scene parsing. Their approach introduces three key contributions: an adaptive weighted downsampling (AWD) layer incorporating spatial and channel information for downsampling, a smooth-oriented convolutional block replacing standard 33 convolutions during training and re-parameterizing back into a 33 convolution block, and a disturbance suppression loss (DSL) measuring Euclidean distance between feature maps from clean and noisy images. Training involves the synthesis of both a clean RAW low-light dataset and a noisy RAW low-light dataset for the disturbance suppression loss. They provided a low-light dataset with paired short-exposure/long-exposure images for instance segmentation training and validation.
3 Synthetic Low-Light Pipeline
For training and evaluating our method, we require annotated low-light images for instance segmentation. However, to the best of our knowledge, there currently exists only one small dataset, released by Chen et al. [11]. This real-world low-light dataset however contains only 2230 paired short-exposure/long-exposure images with instance-level segmentation annotations. Due to the limited number of training paired samples with imagery only in extreme low-light conditions, we decided not to use this dataset. Instead we chose to synthetically create an annotated low-light dataset.
We used the popular Microsoft COCO [1] dataset and passed it through the synthetic low-light pipeline proposed by Lv et al. [14]. This method was chosen over a simplified model (e.g. [25]) because of the simulated in-camera processing which also factors in spatial correlations of noise caused by demosaicing thereby generating more realistic noise characteristics.
4 Proposed Method
4.1 Instance Segmentation
We propose an end-to-end method for instance segmentation in low-light without the need for paired normal-light images. Following the popular Mask R-CNN [5], the network is constructed of a backbone, a Region Proposal Network (RPN), Region of Interest (RoI) pooling, and box and mask heads.
The image is first passed into the modified backbone to extract features. The feature maps are then passed into the RPN first introduced by [15] to generate object proposals with a objectiveness score. These proposals and the feature maps are passed into the RoIPool layer which extracts the feature maps within the proposals. These RoIs are then passed into the box head and mask head to make predictions on the class, bounding box and segmentation mask for each RoI.
For our backbone, we chose to use ResNet-50 [26] with a feature pyramid network (FPN) [27]; this is a popular backbone capable of learning features at different scales. We modify this backbone by adding our weighted NL blocks at the end of res2, res3, res4 and res5 to reduce the feature noise that has been propagated into the feature extractor. The remainder of the network uses the originally proposed parameters from Mask R-CNN [5]. For training, we froze the weights of everything except the backbone to allow the model to specifically learn to denoise the feature maps.
4.2 Weighted non-local blocks for feature denoising
@@
Traditional non-local blocks were first introduced by Wang et al. [13] when they explored implementing non-local means into neural networks. Non-local means [28] is a class of traditional computer vision denoising algorithm that process images on a global scale, unlike local smoothing filters (e.g. bilateral filter [29], Gaussian filter etc.) which only consider neighbouring pixels. Non-local means filtering involves calculating the mean of all pixels in the image, with the weighting determined by the similarity of each pixel to the target one. Wang et al. [13] applied this idea in their neural networks to learn long-range dependencies between pixels spatially and temporally. They also found that the self-attention mechanism [16] is a special form of NL operations in the embedded gaussian form [13].
@@
Xie et al. [30] explored using both non-local means algorithms and local smoothing filtering algorithms for feature denoising to defend against adversarial attacks; attacks that intentionally insert noise into images to confuse systems. They found that although the local denoising operations did improve the robustness of these systems, NL operations were more effective.
@@
In this paper we consider three implementations of NL means for denoising: dot product, Gaussian and embedded Gaussian. The NL blocks consist of an implementation of non-local mean followed by a 11 convolution, added to the block’s input via an identity skip connection. They achieve feature denoising as they pass the feature maps through a denoising operation (i.e. non-local means). We modified these blocks to include learnable weight as shown in Fig. 3. This allows the model to determine the amount of denoising for each ResNet layer. We refer to it as a weighted NL block, which computes the following:
@@
Easy |
Medium |
Difficult |
| Ground truth | Pretrained | Finetuned | Ours |
| Method | LLIE | AP | AP | AP | AP | AP | AP |
| Pretrained | - | 6.9 | 12.4 | 6.9 | 2.3 | 7.4 | 12.4 |
| Pretrained | EnlightenGAN [7] | 5.5 | 10.0 | 5.5 | 1.7 | 6.1 | 10.0 |
| Pretrained | Zero-DCE++ [8] | 5.6 | 10.0 | 5.6 | 1.8 | 6.1 | 9.6 |
| Pretrained | AGLLNet [14] | 6.1 | 11.0 | 6.1 | 1.8 | 7.2 | 10.5 |
| Pretrained | RetinexFormer [6] | 5.7 | 10.3 | 5.7 | 1.8 | 6.5 | 9.9 |
| Finetuned | - | 15.9 | 28.6 | 15.6 | 4.8 | 15.8 | 27.8 |
| Ours (weighted non-local block) | - | 16.9 | 30.7 | 16.6 | 5.6 | 17.2 | 28.9 |
| Original | Ground Truth | Low-Light | Proposed |
| EnlightenGAN [7] | Zero-DCE++ [8] | ||
| AGLLNet [14] | RetinexFormer [6] | ||
5 Experimental Results
@@
For all of our experiments, we trained the models for 10 epochs with an initial learning rate of 0.0005 that decreased by a factor of 10 every 3 epochs, with a batch size of 4. Due to limited computational resources, the number of channels inside the weighted NL block after res2 were divided into 4 instead of 2 as proposed originally by Wang et al. [13]. We used the Microsoft COCO [1] dataset and ran it through the low-light synthetic pipeline. We used 118,287 images for training and 5,000 for evaluation. We did not use their test set as it does not have public annotations available. We used the metrics provided by MS COCO: AP, AP, AP, AP, AP, and AP. AP calculates the mean of the average precisions across multiple IoU thresholds, 0.5 to 0.95, in 0.05 increments. AP and AP calculate the average precision with an IoU threshold of 0.5 and 0.75 respectively. AP, AP, and AP calculate the average precision for small, medium and large objects, which are predefined by [1].
Table 1 shows our results compared to several different methods based around Mask R-CNN [5]. We experimented with one-stage methods and two-stage methods (using LLIE methods [8, 7, 14, 6] for the pre-processing step). We did not run our experiments with a weighted non-local block on res2 due to memory limitations.
5.1 One-stage experiments
@@
Our baseline experiments compare our proposed method with a pretrained Mask R-CNN model [5] and also with finetuning the pretrained model on synthetic low-light training data generated from [1]. Table 1 shows a significant improvement between the fine-tuned model and the pre-trained model. Additionally, there is further enhancement when the weighted NL blocks are integrated, resulting in a +1.0 AP improvement.
@@
Both Table 1 and Fig. 4 illustrate the efficacy of our method in enhancing performance, while also indicating areas for further refinement. Generally, the models exhibit proficiency in segmenting larger objects but encounter challenges with smaller ones, likely attributed to injected noise. The easy example case in Fig. 4 had one large object of a cat and all models were able to accurately generate an instance mask for it. In the medium case, the pretrained Mask R-CNN method failed significantly and falsely labelled many pixels due to the low-light conditions. The finetuned Mask R-CNN performed better although it still suffered with accuracy when segmenting the tennis racket, while our method did not have this issue. In the difficult case, there are small cows in the distance but none of the methods were able to pick them up. These findings are backed by the results from Table 1, particularly the metrics AP, AP and AP. These metrics clearly show that our method improves on the finetuned method for different sized objects, indicating that the weighted NL block successfully removes the noise in the feature extractor.
5.2 Two-stage experiments

### v3

# Multi-Scale Denoising in the Feature Space for Low-Light Instance Segmentation

@@ Abstract
Instance segmentation for low-light imagery remains largely unexplored due to the challenges imposed by such conditions, for example shot noise due to low photon count, color distortions and reduced contrast. In this paper, we propose an end-to-end solution to address this challenging task. Our proposed method implements weighted non-local blocks (wNLB) in the feature extractor. This integration enables an inherent denoising process at the feature level. As a result, our method eliminates the need for aligned ground truth images during training, thus supporting training on real-world low-light datasets. We introduce additional learnable weights at each layer in order to enhance the network’s adaptability to real-world noise characteristics, which affect different feature scales in different ways. Experimental results on several object detectors show that the proposed method outperforms the pre-trained networks with an Average Precision (AP) improvement of at least +7.6, with the introduction of wNLB further enhancing AP by upto +1.3.
Index Terms:
@@
Instance segmentation, low-light, feature denoising, non-local means.I Introduction
@@
Instance segmentation is a crucial task in computer vision because it allows systems to interpret scenes with higher levels of detail, by understanding the position and shape of the objects within them. This can be highly beneficial in scenarios such as autonomous driving, video surveillance, multimedia editing, and agricultural monitoring. Its importance is underscored by numerous recent research contributions in this field and various benchmarks and datasets [1, 2, 3, 4].
@@
While its applications span many sectors, there has been limited research on instance segmentation in low-light conditions. This is because low-light images are heavily affected by noise from a low photon count in darker conditions, making the task much more challenging. A common approach is to first apply a low-light image enhancement (LLIE) method (such as [6, 7, 8, 9]) to these images before passing into existing instance segmentation methods. Specifically, Zero-DCE [7], and RetinexFormer [9] also use low-light object detection benchmarks (i.e. [10, 11]) to demonstrate their enhancing performances. To the best of our knowledge to date, the only known work on low-light instance segmentation is by Chen et al. [12], focusing on leveraging the higher bit-depth in RAW images for better performance.
@@
Our goal is to develop a robust and effective end-to-end method for low-light instance segmentation. This implies eliminating any explicit pre-processing steps in order to simplify the mask generating procedure. However, most models are not designed to handle low-light images; in our experiments, we found that both low-contrast and noise impact the performance of these models, as also confirmed by the study in [13].
In this paper, we propose integrating “plug-and-play” denoising blocks into existing feature extractors, for an end-to-end approach to low-light instance segmentation. Our proposed weighted non-local blocks (wNLB) are an extension to non-local blocks (NLB) [14], where the additional learnable weight ensures the network can learn to ignore appropriate noise characteristics at different feature scales. We trained our framework on a synthetic dataset due to the infancy of machine perception tasks in low-light conditions, focusing on sRGB images as they are more widely used than RAW images and are also applicable to video sequences. We used the popular Microsoft COCO [1] dataset and low-light simulation pipeline proposed by [8]. Our comprehensive experiments on both synthetic and real low-light datasets show our wNLB can improve existing networks performances by upto 1.3AP.
II Related Work
II-A Instance Segmentation
@@
Instance segmentation aims to produce a segmentation mask, bounding box and class label for every object detected in an image. One of the most popular techniques for achieving this is Mask R-CNN [5], introduced by He et al. as an extension to the popular object detection model, Faster R-CNN [15]. With the notable work from [16], there has been a surge in transformer-based methods in recent years. For example, DETR-based [17] instance segmentation methods use CNN-based feature extraction backbones and concatenate positional embeddings before passing into a transformer. However, these often have many parameters and require a lot of computational resources to train effectively. One-stage methods [18, 19] aim to remove the efficiency problem by skipping the region proposal stage found in two-stage methods like Mask R-CNN. With less parameters and less computation required, they are often used in real-time situations, e.g. autonomous driving.
II-B Low-light image enhancement
@@
Traditional LLIE methods treat the problem as two separate tasks: contrast enhancement and denoising. Contrast enhancement methods include gamma correction and adaptive histogram equalizatiom (AHE) [20]; however these lack consideration of illumination factors, leading to cognition approaches inspired by Retinex Theory [21]. Contrast-enhancement of low-light images, however, amplifies noise which necessitates the use of denoising algorithms for further refinement.
@@
Recent advancements in LLIE have explored the use of deep learning. Lore et al. first introduced LLNet [22], an autoencoder trained with synthetic low-light images. Supervised methods, including RetinexNet [23], KiND [24], and AGLLNet [8], require large sets of paired normal/low-light images for training, often generated synthetically. Unsupervised methods gained popularity due to the scarcity of paired normal-light images in low-light datasets. Examples include Zero-DCE (and its extension Zero-DCE++) [7]. RetinexFormer [9], a transformer-based method, is currently the state-of-the-art approach, reformulating the Retinex theory [21] to account for noise as well when decomposing images into reflectance and luminance components.
II-C Low-light scene understanding
Low-light scene understanding is an under-explored field due to the need for high-quality annotations. Datasets such as NightCity [25], BDD100K-night [26], ExDark [10] and [11] are datasets that have been released for several low-light scene understanding tasks, however these do not tackle the instance segmentation problem. As far as we are aware, Chen et al. [12] have released the only low-light instance segmentation dataset to date. However, they only provide 2230 images, which is not suitable for training. Furthermore, their research focused on leveraging RAW data for extreme low-light instance segmentation.
III Proposed Method
@@
Non-local means (NLM) [27] is a class of traditional computer vision denoising algorithms that process images on a global scale, unlike local smoothing filters (e.g. bilateral filter [28], Gaussian filter etc.) which only consider neighbouring pixels. Non-local means filtering involves calculating the mean of all pixels in the image, with the weighting determined by the similarity of each pixel to the target one. Wang et al. [14] applied this idea in their neural networks to learn long-range dependencies between pixels spatially and temporally. In their experiments, they found that the self-attention mechanism [16] is a special form of NL operations in the embedded gaussian form [14].
@@
Xie et al. [29] explored using both non-local means algorithms and local smoothing filtering algorithms for feature denoising to defend against adversarial attacks; attacks that intentionally inject noise into images to confuse computer systems. They found that although the local denoising operations also improved the robustness of these systems, NL operations were more effective in reducing feature noise.
@@
NLBs consist of an implementation of the non-local means algorithm (e.g. dot product, concatenation, Gaussian, embedded Gaussian) followed by a 11 convolution, added to the block’s input via an identity skip connection. They achieve feature denoising as they pass the feature maps through an NLM operation.
Our proposed weighted NLBs extend the NLBs to include a learnable weight as shown in Fig. 3. This allows the model to control the amount of denoising at different scales in the feature extractor. Our wNLBs computes the following:
@@
Method | AP | AP50 | AP75 | APS | APM | APL | |
Mask R-CNN | Pre-trained | 6.9 | 12.4 | 6.9 | 2.3 | 7.4 | 12.4 |
Mask R-CNN | Finetuned | 15.9 | 28.6 | 15.6 | 4.8 | 15.8 | 27.8 |
Mask R-CNN | NLB | 16.6 | 30.3 | 16.4 | 5.4 | 16.7 | 28.1 |
Mask R-CNN | wNLB | 16.9 | 30.7 | 16.6 | 5.6 | 17.2 | 28.9 |
YOLOv8 | Pre-trained | 6.3 | 11.1 | 6.3 | 1.8 | 6.8 | 10.7 |
YOLOv8 | Finetuned | 14.3 | 25.6 | 14.3 | 4.0 | 14.0 | 24.1 |
YOLOv8 | NLB | 22.1 | 37.6 | 22.2 | 7.4 | 23.1 | 36.6 |
YOLOv8 | wNLB | 22.0 | 37.5 | 22.1 | 7.4 | 23.2 | 36.5 |
SOLOv2 | Pre-trained | 8.2 | 14.1 | 8.3 | 2.7 | 8.6 | 14.1 |
SOLOv2 | Finetuned | 15.0 | 26.5 | 14.9 | 3.9 | 15.0 | 26.5 |
SOLOv2 | NLB | 15.8 | 27.9 | 15.6 | 4.1 | 15.8 | 27.6 |
SOLOv2 | wNLB | 15.8 | 27.9 | 15.8 | 4.1 | 15.7 | 27.8 |
Method | AP | AP50 | AP75 | APS | APM | APL |
EnlightenGAN [6] | 5.5 | 10.0 | 5.5 | 1.7 | 6.1 | 10.0 |
ZeroDCE++ [7] | 5.6 | 10.0 | 5.6 | 1.8 | 6.1 | 9.6 |
AGLLNet [8] | 6.1 | 11.0 | 6.1 | 1.8 | 7.2 | 10.5 |
RetinexFormer [9] | 5.7 | 10.3 | 5.7 | 1.8 | 6.5 | 9.9 |
Ours | 16.9 | 30.7 | 16.6 | 5.6 | 17.2 | 28.9 |
IV Experimental Results
For all of our experiments we finetuned the models with an initial learning rate of for 10 epochs. We include the performances from pre-trained models to demonstrate that models trained on popular datasets are not sufficient for low-light. To validate the effectiveness of our proposed wNLB, we used 3 different instance segmentation methods: Mask R-CNN [5], YOLOv8 [18] and SOLOv2 [19]. We used a ResNet-50 FPN [30, 31] backbone for Mask R-CNN and SOLOv2, and a medium-sized CSPDarkNet [18] backkone for YOLOv8 as they are of similar sizes to ensure a fair comparison. Due to limited computational resources (RTX3090), we were unable to add wNLBs at the end of the first layer of the backbones.
@@ IV Experimental Results
We used the Microsoft COCO [1] dataset and passed the training and validation sets through the low-light synthetic pipeline proposed by Lv et al. [8] to create our dataset for training and evaluation. We used their validation set for evaluation as their test set annotations are not publicly available. We also used the metrics provided by MS COCO: AP calculates the mean of the average precisions across multiple IoU thresholds, 0.5 to 0.95, in 0.05 increments. AP50 and AP75 calculate the average precision with an IoU threshold of 0.5 and 0.75 respectively. APS, APM, and APL calculate the average precision for small, medium and large objects, which are pre-defined by [1].
We experimented with one-stage methods (end-to-end frameworks) and two-stage methods (low-light enhancement followed by instance segmentation).
IV-A One-stage experiments
@@
Our baseline experiments compare our proposed method against pre-trained models, both before and after fine-tuning them on synthetic low-light training data as described above. Table I shows a significant improvement between the fine-tuned model and the pre-trained model. Additionally, there is further enhancement when the wNLBs are integrated, resulting in an improvement of at least +0.8 AP.
Fig. 4 visualises the shallow and deep features extracted from the backbones for the pre-trained, finetuned and our proposed methods. The noisy feature maps extracted from the pre-trained method clearly demonstrate that feature noise, caused by low-light conditions, confuse systems leading to poor performance. Upon finetuning the pre-trained model, we find that the feature extractor exhibits an excessive degree of denoising while our wNLBs effectively denoise the signal at multiple scales while preserving essential features, which results in a higher-quality semantic output.
@@
Fig. 5 illustrates the efficacy of our wNLB method in enhancing features, while also indicating areas for further refinement. Generally, the methods exhibit proficiency in segmenting larger objects but encounter challenges with smaller ones, likely attributed to injected noise. The easy example case in Fig. 5 had one large object of a cat and all models were able to accurately generate an instance mask for it. In the medium case, the pre-trained Mask R-CNN method failed significantly and falsely labelled many pixels. The finetuned Mask R-CNN performed better although it still suffered with accuracy when segmenting the tennis racket, while our method did not have this issue. In the difficult case, none of the cows in the distance were detected by the methods.
Despite the marginal gains in AP seen in Table I, we would like to highlight the qualitative improvements, seen in 5, especially for the medium case. We believe that it is important to observe the performance of the methods both quantitatively and qualitatively.
IV-B Two-stage experiments
We also compared our method with the two-stage approaches, where the pre-processing employed popular LLIE methods, EnlightenGAN [6], ZeroDCE++ [7], AGLLNet [8], and RetinexFormer [9] using off-the-shelf weights provided by the authors.